OpenStudy (anonymous):

Find the volume of a cylinder with a height of 4 meters and a diameter that is 3 times the measure of the height. Use 3.14 for pi and round your answer to the nearest hundredth

OpenStudy (midhun.madhu1987):

height, h = 4m diameter, d= 3xh = 3x4 = 12m so raduis r = d/2 = 12/2 = 6m Volume of Cylinder, V = pi x r^2 x h